sir pls pls solve this quotion


Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Rohan

One thing should be noted here that length of the rope is constant.

let the length between the wedge and lower pully is x2

and    length between the block and upper pully is x1

 

           +dx1/dt =v1

and      -dx2/dt =v2

 

Now length of the rope is constant

   so   x2+ (x22+l2)1/2 +x1 =coonstant    where l is a vertical distance between two pully which is constant

now differentiate and put +dx1/dt =v

and      -dx2/dt =v2

and  sinθ =x2/(x22+l2)1/2

and rearrange you will get

v1=v2(1+sinθ)


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best rohan.
 
Regards,
Askiitians Experts
Badiuddin